wrkwhrkhr:

Barbie is analyzing a circle, y2 + x2 = 16, and a linear function g(x). Will they intersect? y2 + x2 = 16 g(x) graph of the function y squared plus x squared equals 16 x g(x) 0 6 1 3 2 0 a Yes, at positive x-coordinates or zero b Yes, at negative x-coordinates or zero c Yes, at negative and positive x-coordinates or zero d No, they will not intersect

wrkwhrkhr:

wrkwhrkhr:

UkuleleGirl:

Ok lets see. What do you know about graphing?

Joe348:

Ok so \[g(x)=-3x+6\]

UkuleleGirl:

hmph joe ur mean I wanted to help but ill let you

Joe348:

@joe348 wrote:
Ok so \[g(x)=-3x+6\]
plug this equation in for Y

Joe348:

@joe348 wrote:
@joe348 wrote:
Ok so \[g(x)=-3x+6\]
plug this equation in for Y
\[(-3x+6)^2+x^2=16\]

Joe348:

@wrkwhrkhr wrote:
Is that it,or is there more to it @joe348 .Overall,it is starting to make sense....
there is more too it

Joe348:

if you still want my help?

Tranquility:

Joe is just making it more complicated than it has to be

Joe348:

true true sorry

Joe348:

you can just calculate it on desmos.com \[y^2+x^2=16\] \[(0,6)\] \[(1,3)\] \[(2,0)\]

Tranquility:

You just need to plot the three points they gave and draw a line through it and see whether or not it intersects the circle. If so, where does it intersect? Is it going to be a positive x-coordinate, negative x-coordinate or at 0?

Joe348:

but you never know if they can't do that traq so I was trying to give them another option.

Tranquility:

You didn't even explain where you got g(x) = -3x + 6 If they can't plot points, they won't have any idea that you're determining the equation for those set of points and then trying to find the point(s) at where the circle and line intersect

Tranquility:

@joe348 wrote:
you can just calculate it on desmos.com \[y^2+x^2=16\] \[(0,6)\] \[(1,3)\] \[(2,0)\]
What are they supposed to calculate here

Joe348:

on a calculator you can just press y= and type the equation

Tranquility:

I know that but that doesn't solve an equation by inputting it

Joe348:

Okay so like on the Y value it decreases by 3 each time til it reaches 0 while x increases by 1 so when i plugged in for x i just add the total value of y and divided by 3 for x

Joe348:

but you seem to got it so i advise you to help them out🤷

Tranquility:

you just switched from explaining how to calculate "it" on desmos to how you determined the equation got me all confused for a second

Tranquility:

@wrkwhrkhr can you graph the points (0, 6) and (1, 3) and (2, 0) on this

1 attachment
Joe348:

@tranquility wrote:
Joe is just making it more complicated than it has to be
.

Joe348:

I was trying to work on the other solution in case they couldn't access that or something

Tranquility:

There's only one solution

Joe348:

I meant other ways to solve for the solution

Tranquility:

@wrkwhrkhr wrote:
Thank you everyone deeply for trying to help in any way,though,can anyone clarify exactly what the overall outcome would be...?Out of a-d.In the end,each explanation helped in so many ways,thank you again.
We're here to help you learn and us giving you the final answer without any effort from your part would defeat the purpose :) If you would take 10 seconds of your time to plot the three points on the graph of the circle, we can get you to the final answer very quickly after that

wrkwhrkhr:

As for @ukulelegirl and @joe348 ,thank you as well for trying to help at your best possible way;I present you both an imaginary medal.As for @tranquility ,thank you for your time and patience,throughout.

Tranquility:

Draw a line between those three points Does that line go through the circle?

Tranquility:

Exactly. You get something like this

1 attachment
wrkwhrkhr:

Yes,totally;thank you so much for the clarification .I just needed a guide to see if I was on track,I just need more confidence in my mathematical skills;thank you again.

Tranquility:

To choose your final answer, you need to say whether it's a positive or negative x-value for where it will intersect Hopefully this will help you (if you haven't already figured it out)

1 attachment
wrkwhrkhr:

Well from my overall knowledge of graphing and such,the slope of the line if falling so would it be negative in outcome...@tranquility ...?

Tranquility:

The line has a negative slope, that's correct. However, we're trying to see at what x-values the line and the circle intersect

Tranquility:

I've drawn an arrow to the two places at where they intersect Is that a positive or negative x-coordinate

Tranquility:

@wrkwhrkhr wrote:
Oh okay,well then I believe that the values of x of the interception between the line & circle is;positive in value...?
That is correct

Tranquility:

(forgot to attach the image in my previous reply but for future reference)

1 attachment
wrkwhrkhr:

Well thank you @tranquility ,it is highly appreciated.Have a good rest of your day.

Tranquility:

It was my pleasure! You too!